# Estimation and goodness-of-fit testing for positive random variables with explicit Laplace transform

### v1

@@ 2 Preliminaries
Let us consider a positive random variable (r.v.) $X$ defined on the probability space $(\Omega,{\mathcal{A}},P)$ and let $L_{X}$ be the Laplace transform of $X$, i.e.
@@
where $\gamma$ is a parameter defined on a suitable subset of ]0,1]\,]0,1]] 0 , 1 ], while $g$ is an appropriate function (see e.g. Barabesi,, 2020, Torricelli et al.,, 2022, and references therein). The family (1) includes some rather exotic distributions, such as the generalized Jacobi laws, which have Laplace transforms given by:
@@
The family with Laplace transform given by (1) encompasses the positive stable distribution. Indeed, a positive stable r.v. has Laplace transform
@@
where $(\gamma,\lambda)\in\,]0,1]\,\times\,]0,\infty[$. For more details on the properties of this r.v. and its stochastic representations, see e.g. Devroye and James, (2014). The integer-valued counterpart of the positive stable distribution is considered in Barabesi and Pratelli, 2014b and Marcheselli et al., (2008). Parameter estimation for the positive stable distribution is discussed in depth in Section 4.
It should be remarked that $g$ could be in turn the Laplace transform of a further positive (not necessarily absolutely-continuous) r.v. $V$. More precisely, if
@@
$L_{X}(s)=E(e^{-sV^{1/\gamma}Z})=E(e^{-s^{\gamma}V})=L_{V}(s^{\gamma}).$
Loosely speaking, in this setting the parameter $\lambda$ is assumed to be a positive r.v. with a suitable distribution. The resulting family encompasses many distributions as the positive Linnik distribution which is obtained when $V$ is distributed according to the Gamma distribution, see e.g. Barabesi et al., 2016a , Huillet, (2000) and Jose et al., (2010).
@@
$L_{X}(s)=g((\theta+s)^{\gamma}-\theta^{\gamma}),$ (4)
where $g$ is an appropriate function and $\theta\geq 0$. If $\theta>0$, the parameter $\gamma$ may also be defined on ],1]\,]$\infty$,1]] - ∞ , 1 ]. Obviously, for $\theta=0$ the family given in (1) is obtained. In turn, the function $g$ in (4) could be a Laplace transform of a further positive r.v. $V$. These distribution classes of such a type are considered by Barabesi et al., 2016a , James, (2010) and Torricelli et al., (2022), among others.
@@
The typical member of the family with Laplace transform given by (4) is the Tweedie distribution. Actually, a Tweedie r.v. has the Laplace transform
$L_{X}(s)=e^{{\rm sgn}(\gamma)\lambda(\theta^{\gamma}-(\theta+s)^{\gamma})},$ (5)
where $(\gamma,\lambda,\theta)\in\{[-\infty,0]\times\,]0,\infty[\times\,]0,\infty[\,\}\cup\{\,]0,1]\times\,]0,\infty[\times[0,\infty[\,\}$. The distribution has been popularized and analysed at length by Hougaard, (1986) and Jørgensen, (1987), following the seminal ideas described in Tweedie et al., (1984). The distribution is central in actuarial studies and ecology, among others (see e.g. Jørgensen,, 1987 and Kendall,, 2004). For its integer-valued counterpart, see Baccini et al., (2016) and Barabesi and Pratelli, 2014a . As to the random variate generation of a Tweedie r.v., see Barabesi and Pratelli, 2014b , Barabesi and Pratelli, (2015), and references therein. The family is very flexible and can be adopted as a model for datasets with complex and non-usual features, since it can simultaneously arrange fat tails and structural zeroes. The inferential issues for the Tweedie law will be discussed in details in Section 5.
@@ 3 Parameter estimation and GOF testing
Let the Laplace transform $L_{X}$ of the target random variable $X$ belong to the general class (4), specifically given by $L_{X}(s)=g\big{(}\lambda((\theta+s)^{\gamma}-\theta^{\gamma})\big{)},$ where $g$ is a suitable infinitely differentiable function on ${\mathbb{R}}\setminus\{0\}$, $\gamma\neq 0$ and $\lambda\kern-0.85358pt>\kern-0.85358pt0$. In particular, $E[X]$ is not finite for $\theta=0$, $\lambda>0$ and $\gamma\in\,]0,1[$. Our objective is to estimate parameters such as $\gamma$, $\lambda$, or $\gamma$, $\lambda$, $\theta$, and to obtain GOF test statistics by empirically estimating $L_{X}$ and its derivatives evaluated only at an ‘appropriate’ random point $A>0$.
Choosing a fixed deterministic point $a>0$ for calculating $L_{X}$ does not provide robust estimations of model parameters. This is because $L_{X}(a)$ could assume negligible values or values practically equal to one for many pairs of parameters $\lambda$ and $\gamma$, with $\theta=0$ or $\theta>0$, making any information on the parameters irrelevant. This drawback persists even when multiple values of $a$ are fixed deterministically and $L_{X}(a)$ or its standard empirical estimates are considered. To bypass this issue, in this section, we introduce a methodology that is broadly applicable to more general classes than (4), enabling robust and informative parameter estimations. In the following sections, we will illustrate these general results in detail for positive stable distributions and Tweedie distributions.
@@
Let us start by noting that for any positive random variable $X$, the following relationship holds:
$L_{X}^{(r)}(s)=(-1)^{r}E(X^{r}e^{-sX}),\qquad\forall s>0,\ \forall r\in\mathbb{N}.$ (6)
In addition, let $T$ be a standard Exponential random variable, independent of $X$. By considering the further random variable $Y_{a}=X\mathbf{1}_{\{aX<T\}},$ where $a>0$ and $\mathbf{1}_{B}$ is the indicator function of the set $B$, it follows that
$L_{Y_{a}}(s)$ $=P(T\leq aX)+L_{X}(s+a)=1-L_{X}(a)+L_{X}(s+a).$
The exponentially censored random variable $Y_{a}$ has a simple expression for its Laplace transform and moments. Even if $E(X)$ is not finite, for $a>0$ we have
$E(Y^{r}_{a})=E(X^{r}e^{-aX})$ (7)
thanks to (6). Thus, $Y_{a}$ may be more manageable than the original random variable $X$, as its moments are available. Equation (7) suggests considering $E(X^{r}e^{-aX})$ for a suitable $a$ when $L_{X}$ depends on some parameter to be estimated.
Remark 1.
If $L_{X}(s)=g\big{(}\lambda((\theta+s)^{\gamma}-\theta^{\gamma})\big{)}$, after some algebra, we also have
$E(Y^{2}_{a})=E(Y_{a})^{2}{{g^{\prime\prime}((\theta+a)^{\gamma}-\theta^{\gamma})}\over{g^{\prime}((\theta+a)^{\gamma}-\theta^{\gamma})^{2}}}+E(Y_{a}){{1-\gamma}\over{\theta+a}},$ (8)
and
$\kern-7.68222ptE(Y^{3}_{a})\kern-1.70717pt=\kern-1.70717ptE(Y_{a})^{3}\,{{g^{\prime\prime\prime}((\theta\kern-1.13809pt+\kern-1.13809pta)^{\gamma}-\theta^{\gamma})}\over{g^{\prime}((\theta\kern-1.13809pt+\kern-1.13809pta)^{\gamma}-\theta^{\gamma})^{3}}}\kern-1.99168pt+\kern-1.99168pt{{1-\gamma}\over{\theta\kern-1.13809pt+\kern-1.13809pta}}\Big{(}3E(Y^{2}_{a})\kern-1.70717pt+\kern-1.70717ptE(Y_{a})\,{{2\gamma\kern-1.13809pt-\kern-1.13809pt1}\over{\theta+a}}\Big{)}.$ (9)
@@
To estimate the model parameters, a natural approach could rely on (6) by adopting the substitution principle based on the empirical counterpart of $E(X^{r}e^{-aX})$, where $r$ does not exceed the number of parameters. In practical applications, the number of parameters typically is $1$, $2$, or $3$, making this method feasible for implementation with real data.
However, the choice of $a$, denoted by $a_{*}$, is crucial for this purpose as discussed earlier in this section. The ‘optimal’ $a_{*}$ will depend on the model parameters, and we propose a data-driven method for obtaining a coherent estimator $A$ of $a_{*}$. To achieve this, first consider that $L_{X}([0,\infty[)=]p,1]$, where $p=P(X=0)$.
@@
If $p=0$, to avoid $Y_{a_{*}}$ being too censored or scarcely censored and thus losing information on parameters, we could set $a_{*}=L_{X}^{-1}(c_{*})$ where $L^{-1}_{X}$ denotes the inverse function of $L_{X}$ and $c_{*}$ is an element of $[1/5,4/5]$. Moreover, in the setting of positive stable distribution, it is convenient to maximize $-c_{*}\log(c_{*})$ for estimating $\gamma$. Since $c_{*}=1/e$ is the maximum point, in the following, we will take
@@
when $p=0$. This implicit definition of $A$ ensures that $A$ is data-driven, making it adaptive to the sample data $X_{1},\ldots,X_{n}$. Since $\cap_{i=1}^{n}\{X_{i}>0\}$ is a negligible event when $p=0$, on the basis of the Glivenko-Cantelli Theorem, $A$ is well defined and converges almost surely to $a_{*}=L^{-1}_{X}(e^{-1})$ because
@@
If $p<1/e$ then the random vector $(\widehat{m}_{1,A},\ldots,\widehat{m}_{k,A},A)$ converges almost surely to the vector $(m_{1,a_{*}},\ldots,m_{k,a_{*}},a_{*})$, while the random vector
$\sqrt{n}(\widehat{m}_{1,A}-m_{1,a_{*}},\ldots,\widehat{m}_{k,A}-m_{k,a_{*}},A-a_{*})$
converges in distribution to the multivariate normal law $\mathrm{N}_{k+1}(0,\Sigma)$ as $n\to\infty$. Moreover, $\Sigma$ is the variance-covariance matrix of the random vector $(V_{1},\ldots,V_{k},W)$, with
@@
In particular, a consistent estimator of $\Sigma$ is given by the sample variance-covariance matrix of the random vectors $(\widehat{V}_{1,i},\ldots,\widehat{V}_{k,i},\widehat{W}_{i})$, $i=1,\ldots,n$, where
@@
$U_{n}=\sqrt{n}\Big{(}\sum_{r=1}^{k}u_{r}(\widehat{m}_{r,a}-m_{r,a})+(A-a)(t-\sum_{r=1}^{k}{u_{r}\over n}\sum_{i=1}^{n}X_{i}^{r+1}e^{-h(A,X_{i})X_{i}})\Big{)}.$
@@
almost surely. Moreover, if $p=0$ there exists a negligible event $H$ such that $s\mapsto L_{n}(s)(\omega)$ is a decreasing function for any $\omega\in H^{c}$. These functions are invertible for any $\omega\in H^{c}$ and their inverse mappings are differentiable at $L_{n}(a)(\omega)$ with derivative $1/L^{\prime}_{n}(a)(\omega)$. Then $\sqrt{n}(A-a)$ coincides almost surely with
$\sqrt{n}(L^{-1}_{n}(e^{-1})-L^{-1}_{n}(L_{n}(a)))\kern-1.70717pt=\kern-1.70717pt{{\sqrt{n}}\over{L^{\prime}_{n}(a)}}(e^{-1}-L_{n}(a))\kern-1.42262pt+\kern-1.42262pto_{P}(\sqrt{n}(e^{-1}-L_{n}(a)))$
Since $L^{\prime}_{n}(a)$ converges almost surely to $-m_{1,a}$, and $o_{P}(\sqrt{n}(e^{-1}-L_{n}(a)))$ coincides with $o_{P}(\sqrt{n}(L_{X}(a)-L_{n}(a)))=\sqrt{n}(L_{X}(a)-L_{n}(a))o_{P}(1),$ from the convergence in distribution of $\sqrt{n}(L_{X}(a)-L_{n}(a))$ it follows that $U_{n}$ has the same large-sample behaviour of the r.v.
$S_{n}=\sqrt{n}\Big{(}\sum_{r=1}^{k}u_{r}(\widehat{m}_{r,a}-m_{r,a})+{{1}\over{m_{1,a}}}(L_{n}(a)-e^{-1})(t-\sum_{r=1}^{k}u_{r}m_{r+1,a})\Big{)}.$
@@
$S_{n}=\,$ $\sqrt{n}\Big{(}\sum_{r=1}^{k}{{u_{r}}\over{n}}\sum_{i=1}^{n}\big{(}(X_{i}^{r}e^{-aX_{i}}-m_{r,a})-{{m_{r+1,a}}\over{m_{1,a}}}(e^{-aX_{i}}-e^{-1})\big{)}$
$+{{t}\over{nm_{1,a}}}\sum_{i=1}^{n}(e^{-aX_{i}}-e^{-1})\Big{)}.$
@@
From Proposition 1, it follows that $\sqrt{n}(A-a_{*})$ converges in distribution to the normal law $\mathrm{N}(0,(L_{X}(2a_{*})-e^{-2})/m_{1,a_{*}}^{2})$ and has the same large-sample behaviour as the r.v. ${{\sqrt{n}}\over{m_{1,a_{*}}}}\big{(}{{1}\over{n}}\sum_{i=1}^{n}e^{-a_{*}X_{i}}-e^{-1}\big{)}.$
@@
Let us now suppose that the distribution of $X$ depends on $k$ parameters, say $\alpha_{1},\ldots,\alpha_{k}$. Let us also suppose that, on the basis of (6), there exist $k$ differentiable functions $h_{1},\ldots,h_{k}$ and a natural number $r\leq k$, such that $\alpha_{j}=h_{j}(m_{1,a_{*}},\ldots,m_{r,a_{*}},a_{*})$ for $j=1,\ldots,k$. Eventually, some $h_{j}$ could be solely functions of $a_{*}$ and in this case we take $r=0$. Thanks to Proposition 1, consistent and large-sample normal estimators ${\widehat{\alpha}_{j}}$ of $\alpha_{j}$, for $j=1,\ldots,k$, are given by
@@
$T_{n}=\sqrt{n}(\widehat{m}_{r+1,A}+(-1)^{r}\widehat{L}_{X}^{(r+1)}(A)),$ (10)
where $\widehat{L}_{X}^{(r)}$ represents the $r$-th derivative of $\widehat{L}_{X}$. Other GOF test statistics can be obtained by substituting $r+1$ with a natural number $q>r+1$. If $h_{1},\ldots,h_{k}$ and $\widehat{L}_{X}$ are functional independent, a further GOF test statistic is given by
$T^{\prime}_{n}=\sqrt{n}(A-\widehat{L}_{X}^{-1}(e^{-1})).$ (11)
@@
$\sqrt{n}\big{(}\psi(\widehat{m}_{1,A},\ldots,\widehat{m}_{r+1,A},A)-\psi(\widehat{m}_{1,a_{*}},\ldots,\widehat{m}_{r+1,a_{*}},a_{*})\big{)}$
where $\psi$ is a differentiable function. From Proposition 1 and the Delta Method, it follows that $T_{n}$ (or $T^{\prime}_{n}$) converges in distribution to a suitable centred normal law. The asymptotic variance of $T_{n}$ (or $T^{\prime}_{n}$) will be explicitly derived for positive stable distribution (or Tweedie distribution) in the following two sections.
@@
A significant, even if simple, illustration of the suggested methodology is obtained by considering a r.v. $X$ with law defined by means of the Laplace transform (2). If $c=\log(e+\sqrt{e^{2}-1})$, from
@@
$T_{n}=\sqrt{n}(\widehat{m}_{1,A}-e^{-2}c\,{\rm sinh}(c)A^{-1}\widehat{\gamma}).$
By using Proposition 1 and the Delta Method, $\sqrt{n}(\widehat{\gamma}-\gamma)$ and $T_{n}$ converge in distribution to suitable centred normal laws, where the variances can be consistently estimated by substituting $\widehat{\gamma}$ and $A$ for $\gamma$ and $a_{*}$ in the corresponding theoretical expressions.
The results for the distribution family considered in the previous example, depending on a single parameter, are appealing and relatively simple. A more complicated setting occurs for a family depending on two or three parameters, as shown in Section 4 and in Section 5, even if the results are rather similar in concept.

### v2

# Estimation and goodness-of-fit testing for non-negative random variables with explicit Laplace transform

@@ 2 Preliminaries
Let us consider a non-negative random variable (r.v.) $X$ defined on the probability space $(\Omega,{\mathcal{A}},P)$ and let $L_{X}$ be the Laplace transform of $X$, i.e.
@@
where $\gamma$ is a parameter defined on a suitable subset of ]0,1]\,]0,1]] 0 , 1 ], while $g$ is an appropriate function (see e.g. Barabesi,, 2020, Torricelli et al.,, 2022, and references therein).
Example 1.
The family (1) includes some rather exotic distributions, such as the generalized Jacobi laws, which have Laplace transforms given by
@@
The family with Laplace transform given by (1) encompasses the positive stable distribution. Indeed, a positive stable r.v. with parameters $(\gamma,\lambda)$, denoted by ${\rm PS}(\gamma,\lambda)$, has Laplace transform
@@
where $(\gamma,\lambda)\in\,]0,1]\,\times\,]0,\infty[$. For more details on the properties of this r.v. and its stochastic representations, see e.g. Devroye and James, (2014). The integer-valued counterpart of the positive stable distribution is considered in Steutel and van Harn, (1979), Barabesi and Pratelli, 2014b and Marcheselli et al., (2008). Parameter estimation for the positive stable distribution is discussed in depth in Section 4.
It should be remarked that $g$ could be in turn the Laplace transform of a further positive (not necessarily absolutely-continuous) r.v. $V$, that is $g=L_{V}$. More precisely, if
@@
$L_{X}(s)=E(e^{-sV^{1/\gamma}Z})=E(e^{-s^{\gamma}V})=L_{V}(s^{\gamma})=g(s^{\gamma}).$
Loosely speaking, in this setting the parameter $\lambda$ is assumed to be a positive r.v. with a suitable distribution. The resulting family encompasses many distributions including the positive Linnik distribution which is obtained when $V$ is distributed according to the Gamma distribution, see e.g. Barabesi et al., 2016a , Huillet, (2000) and Jose et al., (2010).
@@
$L_{X}(s)=g(\{\theta+s\}^{\gamma}-\theta^{\gamma}),$ (4)
where $g$ is an appropriate function and $\theta\geq 0$. If $\theta>0$, the parameter $\gamma$ may also be defined on ],1]\,]$\infty$,1]] - ∞ , 1 ]. Obviously, for $\theta=0$ the family given in (1) is obtained. In turn, the function $g$ in (4) could be a Laplace transform of a further positive r.v. $V$. Distribution classes of such a type are considered by Barabesi et al., 2016a , James, (2010) and Torricelli et al., (2022), among others.
@@
The typical member of the family with Laplace transform given by (4) is the Tweedie distribution. A Tweedie r.v. with parameters $(\gamma,\lambda,\theta)$, denoted by TW$(\gamma,\lambda,\theta)$, has the Laplace transform
$L_{X}(s)=e^{{\rm sgn}(\gamma)\lambda\{\theta^{\gamma}-(\theta+s)^{\gamma}\}},$ (5)
where $(\gamma,\lambda,\theta)\in\{[-\infty,0]\times\,]0,\infty[\times\,]0,\infty[\,\}\cup\{\,]0,1]\times\,]0,\infty[\times[0,\infty[\,\}$. The distribution has been popularized and analysed at length by Hougaard, (1986) and Jørgensen, (1987), following the seminal ideas described in Tweedie et al., (1984). The distribution has been often used in actuarial studies and ecology, among others (see e.g. Jørgensen,, 1987 and Kendall,, 2004). For its integer-valued counterpart, see Baccini et al., (2016) and Barabesi and Pratelli, 2014a . As to the random variate generation of a Tweedie r.v., see Barabesi and Pratelli, 2014b , Barabesi and Pratelli, (2015), and references therein. The family is very flexible and can be adopted as a model for datasets with complex and non-standard features, since it can simultaneously arrange fat tails and structural zeroes. The inferential issues for the Tweedie law will be discussed in details in Section 5.
@@ 3 Parameter estimation and GOF testing
Let the Laplace transform $L_{X}$ of the target random variable $X$ belong to the general class (4), specifically given by $L_{X}(s)=g(\lambda\{(\theta+s)^{\gamma}-\theta^{\gamma}\}),$ where $g$ is a suitable infinitely differentiable function on ${\mathbb{R}}\setminus\{0\}$, $\gamma\neq 0$ and $\lambda\kern-0.85358pt>\kern-0.85358pt0$. In particular, $E(X)$ is not finite for $\theta=0$, $\lambda>0$ and $\gamma\in\,]0,1[$. Our objective is to estimate the parameters and to obtain GOF test statistics by empirically estimating $L_{X}$ and its derivatives at appropriate values.
In this section, we introduce a methodology that is applicable to classes satisfying (4) and more broadly to even more general classes. In the following sections, we will illustrate these general results in detail for positive stable distributions and Tweedie distributions.
@@
Let us start by noting that for any non-negative random variable $X$, the following relationship holds:
$L_{X}^{(r)}(a)=(-1)^{r}E(X^{r}e^{-aX}),\qquad\forall a>0,\ \forall r\in\mathbb{N}.$ (6)
Even if $E(X)$ is not finite, for $a>0$ we have
$E(X^{r}e^{-aX})=E(Y^{r}_{a})$ (7)
@@
where $Y_{a}$ is the exponentially censored random variable $X\mathbf{1}_{\{aX<T\}},$ and $T$ is a standard exponential random variable, independent of $X$. Equation (7) suggests to introduce a variant of the method of moments by considering $E(X^{r}e^{-aX})$ for a suitable $a>0$ when $L_{X}$ depends on some parameters to be estimated. To estimate the model parameters, a natural approach could rely on (6) and on the empirical counterpart of $E(X^{r}e^{-aX})$, where $r$ does not exceed the number of parameters. In practical applications, the number of parameters typically is $1$, $2$, or $3$, making this method feasible for implementation with real data.
Choosing a fixed deterministic point $a>0$ for calculating $L_{X}$ does not provide robust estimations of model parameters. This is because $L_{X}(a)$ could assume negligible values or values practically equal to one for many pairs of parameters $\lambda$ and $\gamma$, with $\theta=0$ or $\theta>0$, making any information on the parameters irrelevant. This drawback persists even when multiple values of $a$ are fixed deterministically, and $L_{X}(a)$ with its empirical counterpart are considered. However, the choice of $a$, denoted by $a_{*}$ is crucial for this purpose. The ‘optimal’ $a_{*}$ will depend on the model parameters, and we propose a data-driven method for obtaining a coherent estimator $A$ of $a_{*}$. To achieve this, first consider that $L_{X}([0,\infty[)=]p,1]$, where $p=P(X=0)$.
@@
If $p=0$, to avoid $Y_{a_{*}}$ being too censored or scarcely censored (i.e. $E(X^{r}e^{-a_{*}X})$ being practically equal to zero or ‘infinite’) and thus loosing information on parameters, we could set $a_{*}=L_{X}^{-1}(c_{*})$ where $L^{-1}_{X}$ denotes the inverse function of $L_{X}$ and $c_{*}$ is an element of $[1/5,4/5]$. Moreover, in the setting of positive stable distribution, it is convenient to maximize $-c_{*}\log(c_{*})$ for estimating $\gamma$. Since $c_{*}=1/e$ is the maximum point, in the following, we will take
@@
when $p=0$. This implicit definition of $A$ ensures that $A$ is data-driven, making it adaptive to the sample data $X_{1},\ldots,X_{n}$. Since $\cup_{i=1}^{n}\{X_{i}=0\}$ is a negligible event when $p=0$, on the basis of the Glivenko-Cantelli Theorem, $A$ is well defined and converges almost surely to $a_{*}=L^{-1}_{X}(e^{-1})$ because
@@
If $p<1/e$ then the random vector $(\widehat{m}_{1,A},\ldots,\widehat{m}_{k,A},A)$ converges almost surely to the vector $(m_{1,a_{*}},\ldots,m_{k,a_{*}},a_{*})$, while the random vector $\sqrt{n}(\widehat{m}_{1,A}-m_{1,a_{*}},\ldots,\widehat{m}_{k,A}-m_{k,a_{*}},A-a_{*})$ converges in distribution to the multivariate normal law $\mathrm{N}_{k+1}(0,\Sigma)$ as $n\to\infty$. Here, $\Sigma$ is the variance-covariance matrix of the random vector $(V_{1},\ldots,V_{k},W)$, with
@@
A consistent estimator of $\Sigma$ is given by the sample variance-covariance matrix of the random vectors $(\widehat{V}_{1,i},\ldots,\widehat{V}_{k,i},\widehat{W}_{i})$, $i=1,\ldots,n$, where
@@
$U_{n}=\sqrt{n}\Big{\{}\sum_{r=1}^{k}u_{r}(\widehat{m}_{r,a}-m_{r,a})+(A-a)(t-\sum_{r=1}^{k}{u_{r}\over n}\sum_{i=1}^{n}X_{i}^{r+1}e^{-h(A,X_{i})X_{i}})\Big{\}}.$
@@
almost surely. Moreover, if $p=0$ there exists a negligible event $H$ such that $s\mapsto L_{n}(s)(\omega)$ is a strictly decreasing function for any $\omega\in H^{c}$. These functions are invertible for any $\omega\in H^{c}$ and their inverse mappings are differentiable at $L_{n}(a)(\omega)$ with derivative $1/L^{\prime}_{n}(a)(\omega)$. Then $\sqrt{n}(A-a)$ coincides almost surely with
$\sqrt{n}\{L^{-1}_{n}(e^{-1})-L^{-1}_{n}(L_{n}(a))\}\kern-1.70717pt=\kern-1.70717pt{{\sqrt{n}}\over{L^{\prime}_{n}(a)}}\{e^{-1}-L_{n}(a)\}\kern-1.42262pt+\kern-1.42262pto_{P}(\sqrt{n}\{e^{-1}-L_{n}(a)\})$
Since $L^{\prime}_{n}(a)$ converges almost surely to $-m_{1,a}$, and $o_{P}(\sqrt{n}\{e^{-1}-L_{n}(a)\})$ coincides with $o_{P}(\sqrt{n}\{L_{X}(a)-L_{n}(a)\})=\sqrt{n}\{L_{X}(a)-L_{n}(a)\}o_{P}(1),$ from the convergence in distribution of $\sqrt{n}\{L_{X}(a)-L_{n}(a)\}$ it follows that $U_{n}$ has the same large-sample behaviour of the r.v.
$S_{n}=\sqrt{n}\Big{[}\sum_{r=1}^{k}u_{r}(\widehat{m}_{r,a}-m_{r,a})+{{1}\over{m_{1,a}}}\{L_{n}(a)-e^{-1}\}(t-\sum_{r=1}^{k}u_{r}m_{r+1,a})\Big{]}.$
@@
$S_{n}=\,$ $\sqrt{n}\Big{[}\sum_{r=1}^{k}{{u_{r}}\over{n}}\sum_{i=1}^{n}\big{\{}(X_{i}^{r}e^{-aX_{i}}-m_{r,a})-{{m_{r+1,a}}\over{m_{1,a}}}(e^{-aX_{i}}-e^{-1})\big{\}}$
$+{{t}\over{nm_{1,a}}}\sum_{i=1}^{n}(e^{-aX_{i}}-e^{-1})\Big{]}.$
@@
From Proposition 1, it follows that $\sqrt{n}(A-a_{*})$ converges in distribution to the normal law $\mathrm{N}(0,\{L_{X}(2a_{*})-e^{-2}\}/m_{1,a_{*}}^{2})$ and has the same large-sample behaviour as the r.v. ${{\sqrt{n}}\over{m_{1,a_{*}}}}\big{(}{{1}\over{n}}\sum_{i=1}^{n}e^{-a_{*}X_{i}}-e^{-1}\big{)}.$
@@
Let us now suppose that the distribution of $X$ depends on $k$ parameters, say $\alpha_{1},\ldots,\alpha_{k}$. Let us also suppose that, on the basis of (6), there exist $k$ differentiable functions $h_{1},\ldots,h_{k}$ and a natural number $r\leq k$, such that $\alpha_{j}=h_{j}(m_{1,a_{*}},\ldots,m_{r,a_{*}},a_{*})$ for $j=1,\ldots,k$. Some $h_{j}$ could be solely functions of $a_{*}$ and in this case we take $r=0$. Thanks to Proposition 1, consistent and large-sample normal estimators ${\widehat{\alpha}_{j}}$ of $\alpha_{j}$, for $j=1,\ldots,k$, are given by
@@
$T_{n}=\sqrt{n}\{\widehat{m}_{r+1,A}+(-1)^{r}\widehat{L}_{X}^{(r+1)}(A)\},$ (8)
where $\widehat{L}_{X}^{(r)}$ represents the $r$-th derivative of $\widehat{L}_{X}$. Other GOF test statistics can be obtained by substituting $r+1$ with a natural number $q>r+1$. If $\widehat{L}_{X}(A)$ is not a degenerate random variable, a further GOF test statistic is given by
$T^{\prime}_{n}=\sqrt{n}\{e^{-1}-\widehat{L}_{X}(A)\}.$ (9)
@@
$\sqrt{n}\big{\{}\psi(\widehat{m}_{1,A},\ldots,\widehat{m}_{r+1,A},A)-\psi(m_{1,a_{*}},\ldots,m_{r+1,a_{*}},a_{*})\big{\}}$
where $\psi$ is a differentiable function. From Proposition 1 and the Delta Method, it follows that $T_{n}$ (or $T^{\prime}_{n}$) converges in distribution to a centered normal law. The asymptotic variance of $T_{n}$ (or $T^{\prime}_{n}$) will be explicitly derived for positive stable distribution (or Tweedie distribution) in the following two sections.
@@
An interesting, even if simple, illustration of the suggested methodology is obtained by considering a r.v. $X$ with law defined by means of the Laplace transform (2). If $c=\log(e+\sqrt{e^{2}-1})$, from
@@
$T_{n}=\sqrt{n}\{\widehat{m}_{1,A}-e^{-2}c\,{\rm sinh}(c)A^{-1}\widehat{\gamma}\}.$
By using Proposition 1 and the Delta Method, $\sqrt{n}(\widehat{\gamma}-\gamma)$ and $T_{n}$ converge in distribution to centered normal laws, where the variances can be consistently estimated by substituting $\widehat{\gamma}$ and $A$ for $\gamma$ and $a_{*}$ in the corresponding theoretical expressions.
The results for the distribution family considered in the previous example, depending on a single parameter, are appealing and relatively simple. A more complicated setting occurs for a family depending on two or three parameters, as shown in Section 4 and in Section 5, even if the results are rather similar in concept. In this more complex framework the following result can be useful.
Remark 4.
If $L_{X}(s)=g(\lambda\{(\theta+s)^{\gamma}-\theta^{\gamma}\})$, after some algebra, we also have
$E(Y^{2}_{a})=E(Y_{a})^{2}{{g^{\prime\prime}(\lambda\{(\theta+a)^{\gamma}-\theta^{\gamma}\})}\over{g^{\prime}(\lambda\{(\theta+a)^{\gamma}-\theta^{\gamma}\})^{2}}}+E(Y_{a}){{1-\gamma}\over{\theta+a}},$ (10)
and
$\kern-7.68222ptE(Y^{3}_{a})\kern-2.56073pt=\kern-2.56073pt{{E(Y_{a})^{3}g^{\prime\prime\prime}(\lambda\{(\theta\kern-1.70717pt+\kern-1.70717pta)^{\gamma}\kern-1.70717pt-\kern-1.42262pt\theta^{\gamma}\})}\over{g^{\prime}(\lambda\{(\theta\kern-1.70717pt+\kern-1.70717pta)^{\gamma}\kern-1.70717pt-\kern-1.42262pt\theta^{\gamma}\})^{3}}}\kern-1.99168pt+\kern-1.99168pt{{1\kern-1.70717pt-\kern-1.70717pt\gamma}\over{\theta\kern-1.13809pt+\kern-1.13809pta}}\Big{\{}3E(Y^{2}_{a})\kern-1.70717pt+\kern-1.70717ptE(Y_{a})\,{{2\gamma\kern-1.13809pt-\kern-1.13809pt1}\over{\theta+a}}\Big{\}}.$ (11)